OpenStudy (anonymous):

Solve using substitution or elemination methods...can someone help me im soo cunfused

OpenStudy (anonymous):

2x+5y=-5 x+3y=3

OpenStudy (anonymous):

Multiply the bottom equation times -2, just make sure you do it to all terms of the equation, including on the other side of the equals sign. Then you can add to equation 1. You should then have one equation with one unknown variable.

OpenStudy (anonymous):

is this the correct work? x=(-5-5y)/2 (-5+y)/2+3y=3 then 6y/6=6/6 so then y=1? and x=(-5-5y)/2=-5?

OpenStudy (anonymous):